What is the volume (in dL) of a box that is 3.7 ft long, 0.532 m wide and 1245 mm tall?
1
Expert's answer
2020-03-04T07:59:27-0500

As 1 ft = 3.048 dm, then the length is ="3.7\times3.048 = 11.28 dm"

As 1 m = 10 dm , then the width is ="0.532\times10=5.32dm"

As 1mm = 0.01 dm, the the hegth is = "1245\times0.01=12.45dm"


"Volume = length\times width\times height = 11.28\times 5.32\times 12.45 = 747 dm^3"


As "1 dm^3 =1 dL"

then the volume of a box is "747 \times 10 = 7470 dL"